# Maximal ideals of reduced group C*-algebras and Thompson’s groups

### v2

@@ 1. Introduction
Recently, Christensen and Neshveyev proved in [CN23, Theorem A] that, given an étale groupoid $\mathcal{G}$, any proper ideal of $C^{*}_{r}(\mathcal{G})$ is contained in a proper ideal induced from an isotropy group of $\mathcal{G}$.
@@
Let $F\leq T$ be Thompson’s groups. It is a long-standing open problem whether $F$ is amenable (see [CFP96] for an introduction to Thompson’s groups). Haagerup and Olesen showed in [HO17, Theorem 5.5] that, if $F$ is amenable, then $C^{*}_{r}(T)$ is not simple. Furthermore, in [LBMB18, Corollary 4.2], Le Boudec and Matte Bon showed the converse. In particular, if $F$ is not amenable, then $C^{*}_{r}(T)$ has a unique maximal ideal (namely, the zero ideal). By using Theorem B and a description due to Dudko and Medynets ([DM14]) of extreme characters on Thompson’s groups, we prove the following:
@@
The reduced C*-algebra of Thompson’s group $T$ has a unique maximal ideal.
@@
These results yield the following characterizations of amenability of $F$ in terms of the ideal structure of $C^{*}_{r}(T)$:
@@ 2. Induction of ideals via conditional expectations
Remark 2.5.
We do not know any example in which $P(J)$ is not closed. In some cases, this is known to hold automatically (see [BFPR, Proposition 3.22] and [CN23, Lemma 2.1]).
@@
The next result says that the pair $(\textup{Coind}_{P},\textup{Ind}_{P})$ forms a Galois connection between $\mathbb{I}(B)$ and $\mathbb{I}(A)$. The result is a consequence of Lemma 2.7 and [DP02, Lemma 7.26].
@@
Let $\textup{Max}(A)$ and $\textup{Max}(B)$ denote the maximal (proper) ideals of $A$ and $B$, respectively. Let $\textup{Max}_{P}(A)$ denote the collection of maximal co-induced (proper) ideals of $A$.
@@
Since $A$ is unital, the join of a family of proper ideals of $A$ is a proper ideal. The result then follows from Zorn’s lemma and the fact that, by part 3 of Proposition 2.9, $\textup{Coind}_{P}$ preserves joins. ∎
@@
Then 1$\iff$2. If $A$ is unital, then 2$\implies$3$\implies$4. If $B$ is unital, then 4$\implies$1.
@@
1$\implies$2. Let $I\coloneqq\textup{Coind}_{P}(J)$. Then $J\subseteq\textup{Ind}_{P}(I)$ by Lemma 2.7.
2$\implies$1 follows from the fact that, by Lemma 2.7,
@@
2$\implies$3. Let $J\in\textup{Max}(B)$. Then, 2 gives us a proper ideal $I\in\mathbb{I}(A)$ such that $J\subseteq\textup{Ind}_{P}(I)$. Since $A$ is unital, we can assume $I$ is maximal. From part 7 of Proposition 2.9, $\textup{Ind}_{P}(I)$ is a proper ideal, so that $J=\textup{Ind}_{P}(I)$ by maximality of $J$.
3$\implies$4. By part 4 of Proposition 2.9, we have that $\textup{Ind}_{P}$ is injective on $\textup{Max}_{P}(A)$. Given $J\in\textup{Max}(B)$, we need to find $M\in\textup{Max}_{P}(A)$ such that $\textup{Ind}_{P}(M)=J$. By assumption, there exists a proper ideal $I\in\mathbb{I}(A)$ such that $J=\textup{Ind}_{P}(I)$. By part 1 of Proposition 2.9, we can replace $I$ with $I^{\prime}\coloneqq\textup{Coind}_{P}(\textup{Ind}_{P}(I))$, which is co-induced and satisfies $J=\textup{Ind}_{P}(I^{\prime})$. By Lemma 2.10, there exists $M\in\textup{Max}_{P}(A)$ such that $I^{\prime}\subseteq M$. Now we have $J=\textup{Ind}_{P}(I^{\prime})\subseteq\textup{Ind}_{P}(M)$. Hence $J=\textup{Ind}_{P}(M)$ by maximality of $J$.
4$\implies$1. Take $M\in\textup{Max}(B)$ such that $J\subseteq M$, and let $I\in\textup{Max}_{P}(A)$ be such that $M=\textup{Ind}_{P}(I)$. Since $\textup{Coind}_{P}(J)\subseteq\textup{Coind}_{P}(\textup{Ind}_{P}(I))\subseteq I$, we conclude that $\textup{Coind}_{P}(J)$ is proper. ∎
@@ 3. Reduced group C*-algebras
We denote by $\textup{Max}_{H}(C^{*}_{r}(H))$ the set of maximal co-induced (proper) ideals of $C^{*}_{r}(H)$.
@@
This is a direct consequence of Theorem 2.11 and [BKKO17, Lemma 7.9]. ∎
@@
=$=\begin{cases}(\lambda_{g}a\lambda_{g}^{*})\lambda_{gh}&\text{ if}gh\in N,\\ 0&\text{ if }gh\not\in N.\end{cases}$
@@
and let $\phi\colon C(X)\rtimes_{r}G_{x}\to C^{*}_{r}(G_{x})$ be the ***-homomorphism given by $\phi(f\lambda_{g})=f(x)\lambda_{g}$, for all $f\in C(X)$ and $g\in G$. Then, $\sigma\coloneqq\tau\circ\phi\circ\psi$ satisfies (3.1).
@@ 4. Application: Thompson’s groups
Let $\mathbb{Z}[1/2]$ be the set of dyadic rationals and let $\Omega\coloneqq[0,1)\cap\mathbb{Z}[1/2]$. Thompson’s group $T$ is the group of homeomorphisms on $\frac{\mathbb{R}}{\mathbb{Z}}$ that are piecewise linear, with finitely many breakpoints, all at $\Omega$, and slopes in $2^{\mathbb{Z}}$. Thompson’s group $F$ is the subgroup of $T$ which stabilizes 00. The commutator $[F,F]$ coincides with the set of elements of $T$ which fix pointwise a neighborhood of 00 ([CFP96, Theorem 4.1]).
@@
Take sequences of dyadic rationals $(y_{n}),(z_{n})\subseteq[0,1)$ such that $y_{n}\uparrow x$ and $z_{n}\downarrow x$. Let $T_{n}$ be the subgroup of $T$ consisting of elements which pointwise fix a neighborhood of the interval $[y_{n},z_{n}]$.
@@
Recall that a tracial state on a unital $C^{*}$-algebra is extremal if and only if its GNS representation is factorial (see the proof of [BdlH20, Proposition 11.C.3], or [HM08, Proposition 1]). In particular, if $\pi\colon B\to A$ is a surjective ***-homomorphism between unital $C^{*}$-algebras, and $\tau\in\partial T(A)$, then $\tau\circ\pi\in\partial T(B)$.
@@
Given a unitary representation $\pi$ of a group $G$, let $C^{*}_{\pi}(G)$ denote the C*-algebra generated by the image of $\pi$. Recall that, if $\sigma$ is another unitary representation of $G$, then $\pi$ is weakly contained in $\sigma$ (denoted by $\pi\prec\sigma$), if there is a (necessarily surjective) ***-homomorphism $C^{*}_{\sigma}(G)\to C^{*}_{\pi}(G)$ which, for each $g\in G$, maps $\sigma_{g}$ to $\pi_{g}$.
@@
In particular, there cannot be a ***-homomorphism $C^{*}_{\lambda_{T/F_{m}}}(T)\to C^{*}_{\lambda_{T/F_{n}}}(T)$ which, for each $t\in T$, maps $\lambda_{T/F_{m}}(t)$ to $\lambda_{T/F_{n}}(t)$. This shows the claim. ∎

### v3

@@ 1. Introduction
Recently, Christensen and Neshveyev proved in [CN24, Theorem A] that, given an étale groupoid $\mathcal{G}$, any proper ideal of $C^{*}_{r}(\mathcal{G})$ is contained in a proper ideal induced from an isotropy group of $\mathcal{G}$.
@@
Since the isotropy groups $G_{x}$ are amenable, Theorem B has the following surprising consequence: it reduces the problem of understanding maximal ideals in reduced group C*-algebras of arbitrary groups to understanding ideals in group C*-algebras of amenable groups.
Given a countable group $G$, Le Boudec and Matte Bon proved in [LBMB18, Section 2.4] that $\mathcal{A}_{G}:=\{G_{x}:x\in\partial_{F}G\}$ is the largest amenable uniformly recurrent subgroup of $G$, and used this fact for computing $\mathcal{A}_{G}$ for several examples of groups. Thus, Theorem B can also be seen as the version for ideals of the fact shown by Breuillard, Kalantar, Kennedy and Ozawa in [BKKO17, Theorem 4.1] that any tracial state on $C^{*}_{r}(G)$ is supported on the largest amenable normal subgroup of $G$ .
@@
Let $F\leq T$ be Thompson’s groups. It is a long-standing open problem to determine whether or not $F$ is amenable (see [CFP96] for an introduction to Thompson’s groups). Haagerup and Olesen showed in [HO17, Theorem 5.5] that, if $F$ is amenable, then $C^{*}_{r}(T)$ is not simple. Furthermore, in [LBMB18, Corollary 4.2], Le Boudec and Matte Bon showed the converse. In particular, if $F$ is not amenable, then $C^{*}_{r}(T)$ has a unique maximal ideal (namely, the zero ideal). By using Theorem B and a description due to Dudko and Medynets ([DM14]) of extreme characters on Thompson’s groups, we prove the following:
@@
The reduced C*-algebra of Thompson’s group $T$ always has a unique maximal ideal.
@@
These results together yield the following characterizations of amenability of $F$ in terms of the ideal structure of $C^{*}_{r}(T)$:
@@
Acknowledgement
We thank Narutaka Ozawa for allowing us to include Example 2.5.
@@ 2. Induction of ideals via conditional expectations
In some cases, $P(J)$ is known to be automatically closed (see [BFPR, Proposition 3.22] and [CN24, Lemma 2.1]). In general, this is false, as the following example, due to N. Ozawa (personal communication), shows.
Example 2.5 (Ozawa).
Take a sequence $(A_{i})_{i\in\mathbb{N}}$ of disjoint finite nonempty subsets of $\mathbb{N}$ such that $|A_{i}|\to+\infty$ and $\mathbb{N}=\bigsqcup_{i\in\mathbb{N}}A_{i}$.
[…]
$P(f)(x):=\sum_{y\in A_{i}}\frac{f(y)}{|A_{i}|},$
for $f\in\ell^{\infty}(\mathbb{N})$, $i\in\mathbb{N}$, and $x\in A_{i}$. Then $P$ is a conditional expectation onto $D:=\{f\in\ell^{\infty}(\mathbb{N}):\text{$\forall i$, $f|_{A_{i}}$ is constant}\}$.
Fix a sequence $(a_{i})_{i\in\mathbb{N}}\subset\mathbb{N}$ such that, for each $i\in\mathbb{N}$, $a_{i}\in A_{i}$. Let $J:=\ell^{\infty}(\bigsqcup_{i\in\mathbb{N}}\{a_{i}\})\unlhd\ell^{\infty}(\mathbb{N})$.
Given $f\in J$ and $i\in\mathbb{N}$, we have that
$P(f)(a_{i})=\frac{f(a_{i})}{|A_{i}|}.$
Therefore, $\overline{P(J)}=\{g\in D:g(a_{i})\to 0\}$. Let $g\in D$ be such that $g(a_{i})=|A_{i}|^{-1/2}$, for all $i\in\mathbb{N}$. Then $g\in\overline{P(J)}\setminus P(J)$.
@@
The next result says that the pair $(\textup{Coind}_{P},\textup{Ind}_{P})$ forms a Galois connection between $\mathbb{I}(B)$ and $\mathbb{I}(A)$. The result is an immediate consequence of Lemma 2.7 (see also [DP02, Lemma 7.26]).
@@
Let $\textup{Max}(A)$ and $\textup{Max}(B)$ denote the maximal (proper) ideals of $A$ and $B$, respectively. Let $\textup{Max}_{P}(A)$ denote the collection of maximal co-induced (proper) ideals of $A$ via $P\colon B\rightarrow A$.
@@
Since $A$ is unital, the join of a chain of proper ideals of $A$ is a proper ideal. The result then follows from Zorn’s lemma and the fact that, by part 3 of Proposition 2.9, $\textup{Coind}_{P}$ preserves joins. ∎
@@
[…]
@@
[…]
@@
[…]
@@ 3. Reduced group C*-algebras
We denote by $\textup{Max}_{H}(C^{*}_{r}(H))$ the set of maximal co-induced (proper) ideals of $C^{*}_{r}(H)$ via $P_{H}$.
The next result is an immediate consequence of [BKKO17, Lemma 7.9] and Theorem 2.11. For completeness, we present a more direct proof. We will use operator algebraic properties of the Furstenberg boundary that can be found in, for example, [BKKO17, Section 2.4]. Given a group $G$ and unital $C^{*}$-algebras $A$ and $B$ endowed with $G$-actions, we call a $G$-equivariant unital completely positive linear map from $A$ to $B$ a $G$-map.
@@
1. Let $J$ be a maximal ideal of $C^{*}_{r}(G)$ and fix $x\in\partial_{F}G$. We claim that there exists a state $\sigma$ on $C^{*}_{r}(G)$ such that $\sigma|_{J}=0$ and, for all $g\in G\setminus G_{x}$, $\sigma(\lambda_{g})=0$.
By $G$-injectivity of $C(\partial_{F}G)$, there is a $G$-map $\varphi\colon\frac{C^{*}_{r}(G)}{J}\to C(\partial_{F}G)$. Let $\pi\colon C^{*}_{r}(G)\to\frac{C^{*}_{r}(G)}{J}$ be the quotient. By $G$-injectivity again, there is a $G$-map $\psi\colon C(\partial_{F}G)\rtimes_{r}G\to C(\partial_{F}G)$ such that $\psi|_{C^{*}_{r}(G)}=\varphi\circ\pi$. By $G$-rigidity, we have that $\psi|_{C(\partial_{F}G)}=\mathrm{Id}_{C(\partial_{F}G)}$. Let $\mathrm{Ev}_{x}\colon C(\partial_{F}G)\to\mathbb{C}$ be evaluation in $x$ and $\sigma:=\mathrm{Ev}_{x}\circ\varphi\circ\pi$. Clearly, $\sigma|_{J}=0$. Given $g\in G\setminus G_{x}$, take $f\in C(\partial_{F}G)$ such that $f(gx)=1$ and $f(x)=0$. Then
$\sigma(\lambda_{g})=\sigma(\lambda_{g})f(gx)=\mathrm{Ev_{x}}(\psi(\lambda_{g}))\mathrm{Ev_{x}}(g^{-1}f)$ $=\mathrm{Ev_{x}}(\psi(\lambda_{g})g^{-1}f)$
$=\mathrm{Ev}_{x}(\psi(f\lambda_{g}))$
$=f(x)\sigma(\lambda_{g})$
$=0,$
thus showing the claim.
Let $K:=\{a\in C^{*}(G_{x}):\sigma(bac)=0\text{ for all $b,c\in C^{*}(G_{x})$}\}$. Notice that $\sigma|_{C^{*}(G_{x})}\circ P_{G_{x}}=\sigma$ and $C^{*}(G_{x})$ is contained in the multiplicative domain of $P_{G_{x}}$. Given $a\in J$, $b,c\in C^{*}(G_{x})$, and $d,e\in C^{*}_{r}(G)$, we have that
$\sigma(bP_{G_{x}}(dae)c)=\sigma(P_{G_{x}}(bdaec))=\sigma(bdaec)=0,$
where the last equality follows from the fact that $\sigma|_{J}=0$. Therefore, $J\subset\textup{Ind}_{G_{x}}(K)$ (hence, by maximality, $J=\textup{Ind}_{G_{x}}(K)$). By taking a maximal ideal $I$ of $C^{*}(G_{x})$ which contains $K$, we obtain that $J=\textup{Ind}_{G_{x}}(I)$.
2 follows from 1 and Theorem 2.11. ∎
Remark 3.2.
Given $x\in\partial_{F}G$, consider the canonical conditional expectation $Q\colon C(\partial_{F}G)\rtimes_{r}G\to C^{*}(G_{x})$. It follows from [CN24, Theorem 3.4] that induction gives a bijection between $\textup{Max}_{Q}(C^{*}(G_{x}))$ and $\textup{Max}(C(\partial_{F}G)\rtimes_{r}G)$. Furthermore, since $Q|_{C^{*}_{r}(G)}=P_{G_{x}}$ and $Q|_{C(\partial_{F}G)}$ is multiplicative, it is not difficult to check that, for any ideal $J$ of $C^{*}(G_{x})$, it holds that $\textup{Ind}_{Q}(J)\cap C^{*}_{r}(G)=\textup{Ind}_{G_{x}}(J)$. Unfortunately, it is not clear whether this line of thinking could lead to a new proof of Theorem 3.1. We thank a referee for suggesting us to include this observation.
@@
$=\begin{cases}(\lambda_{g}a\lambda_{g}^{*})\lambda_{gh}&\text{ if }gh\in N,\\ 0&\text{ if }gh\not\in N.\end{cases}$
@@
and let $\phi\colon C(X)\rtimes_{r}G_{x}\to C^{*}_{r}(G_{x})$ be the *-homomorphism given by $\phi(f\lambda_{g})=f(x)\lambda_{g}$, for all $f\in C(X)$ and $g\in G$. Then, $\sigma\coloneqq\tau\circ\phi\circ\psi$ satisfies (3.1).
@@ 4. Application: Thompson’s groups
Let $\mathbb{Z}[1/2]$ be the set of dyadic rationals and let $\Omega\coloneqq[0,1)\cap\mathbb{Z}[1/2]$. Thompson’s group $T$ is the group of homeomorphisms on $\frac{\mathbb{R}}{\mathbb{Z}}$ that are piecewise linear, with finitely many breakpoints, all at $\Omega$, and slopes in $2^{\mathbb{Z}}$. Thompson’s group $F$ is the subgroup of $T$ which stabilizes 0. The commutator $[F,F]$ coincides with the set of elements of $T$ which fix pointwise a neighborhood of 0 ([CFP96, Theorem 4.1]).
@@
Take sequences of dyadic rationals $(y_{n}),(z_{n})\subseteq[0,1)$ such that $y_{n}\uparrow x$ and $z_{n}\downarrow x$. Let $T_{n}$ be the subgroup of $T$ consisting of elements which pointwise fix a neighborhood of the interval $y_{n},z_{n}]$.
@@
Recall that a tracial state on a unital $C^{*}$-algebra is extremal if and only if its GNS representation is factorial (see the proof of [BdlH20, Proposition 11.C.3], or [HM08, Proposition 1]). In particular, if $\pi\colon B\to A$ is a surjective *-homomorphism between unital $C^{*}$-algebras, and $\tau\in\partial T(A)$, then $\tau\circ\pi\in\partial T(B)$.
@@
Given a unitary representation $\pi$ of a group $G$, let $C^{*}_{\pi}(G)$ denote the C*-algebra generated by the image of $\pi$. Recall that, if $\sigma$ is another unitary representation of $G$, then $\pi$ is weakly contained in $\sigma$ (denoted by $\pi\prec\sigma$), if there is a (necessarily surjective) *-homomorphism $C^{*}_{\sigma}(G)\to C^{*}_{\pi}(G)$ which, for each $g\in G$, maps $\sigma_{g}$ to $\pi_{g}$.
@@
In particular, there cannot be a *-homomorphism $C^{*}_{\lambda_{T/F_{m}}}(T)\to C^{*}_{\lambda_{T/F_{n}}}(T)$ which, for each $t\in T$, maps $\lambda_{T/F_{m}}(t)$ to $\lambda_{T/F_{n}}(t)$. This shows the claim. ∎